OpenStudy (anonymous):

i'm trying to find the centroid of the region bonded by the given curves y= sinx; y=cosx, x=o and x=pi/4

OpenStudy (anonymous):

The first thing you need is the mass so if we set f(x) = sin(x) and g(x) = cos(x) to get the mass we want: \[\int\limits_{0}^{π/4}[\cos(x)-\sin(x)]dx\]

OpenStudy (anonymous):

i know thats part i think im getting the wrong numbers when i plug the integral

OpenStudy (anonymous):

This integral should be equal to √(2) -1

OpenStudy (anonymous):

is that my A?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oh i got 2/ sqrt(2)-1

OpenStudy (anonymous):

but how do you do the integrarion by parts?

OpenStudy (anonymous):

This first integral isn't by parts. ∫cos(x)dx = sin(x) ∫sin(x)dx = -cos(x) sin(x)+cos(x) evaluated from 0 to π/4. At π/4 we get √2 and at 0 we get 1.

OpenStudy (anonymous):

Next we have to find the x and y coordinate.

OpenStudy (anonymous):

For the x coordinate take: \[\int\limits_{0}^{π/4}x[\cos(x)-\sin(x)]dx\] and divide by √2 - 1 and the y coordinate take: \[\int\limits\limits_{0}^{π/4}(1/2)[\cos ^{2}(x) - \sin ^{2}(x)]dx\] and divide that by √2 - 1